Circular No. 7125

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999ap, 1999aq, 1999ar IN ANONYMOUS GALAXIES

L. G. Strolger, University of Michigan; L. C. Ho, Carnegie Observatories; R. Covarrubias, Cerro Tololo Interamerican Observatory (CTIO); M. Hamuy, Steward Observatory; R. C. Smith, CTIO; and the rest of the Nearby Galaxies Supernova Search team, along with G. Aldering, R. Knop, T. York, and M. Kim (Lawrence Berkeley National Laboratory) on behalf of the Supernova Cosmology Project, report the discovery of three supernovae in R-band images taken with the Kitt Peak 0.9-m telescope (+ MOSAIC imager) during their searches on Feb. 21–25 and Mar. 5–9 UT.

SN  1999 UT  α_{2000}  δ_{2000}  R
1999ap Feb. 20 8^{h}29^{m}47.^{s}46 + 4^{o}37′38′′.5 19.4
1999aq Feb. 20 9 35 40.48 − 4 55 23.2 18.8
1999ar Mar. 6 9 20 16.00 + 0 33 39.6 19.7

SN 1999ap appears to be a type-II supernova at redshift 0.04, exhibiting a well-developed Hα P-Cyg profile, based on a spectrum (range 400– 800 nm) obtained on Mar. 5.3 by G. Aldering with the MDM 2.4-m telescope. A spectrum (range 400–1000 nm) of SN 1999aq taken by A. V. Filippenko and W. D. Li (University of California at Berkeley) on Mar. 9 with the Keck II telescope shows that the supernova is of type Ia at a redshift of roughly 0.05, observed ∼ 1 month past maximum brightness. SN 1999ar is confirmed to be a type-Ia event, observed near maximum brightness, at a redshift of 0.15 based on a spectrum (range 320–890 nm; total exposure 4500 s) obtained by B. Schmidt (Mount Stromlo and Siding Spring Observatories) on Mar. 10.6 with the ANU 2.3-m telescope (+ Double Beam Spectrograph). Finding charts for the three objects (identified as NB-102, NB-175, and NB-187, respectively) can be obtained from

http://panisse.lbl.gov/nearsearch/fcharts/index.html.

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates: Jan. 18.45 UT, 10.5 (D. A. J. Sear-

gent, The Entrance, N.S.W., 25×100 binoculars); Feb. 4.44, 11.0 (Seargent, Cowra, N.S.W., 0.25-m reflector); 13.40, 11.1 (S. T. Rae, Te Awamutu, New Zealand, 0.11-m reflector); Mar. 9.53, 11.7 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 14.50, 11.6 (M. Mattiazzo, Wallaroo, South Australia, 0.20-m reflector).

1999 March 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green